Form the difference equation generated by yx=a2x+b3x+c. (5 marks)
1
Expert's answer
2020-06-03T18:46:22-0400

Given "y(x)=a^2 x+b^3x+c = (a^2+b^3) x +c" where "a,b,c" are the parameter.

By differentiation with respect to "x", we get "y' = a^2 + b^3" .

Again differentiation with respect to "x", we get "y'' = 0" which is required differential generated by "y" .